mathslover (mathslover):

Let \(f(x) = \int_{2}^{x} \frac{dt}{\sqrt{1+t^4}} \) and \(g\) be the the inverse of \(f\). Then, the value of \(g'(0)\) is ... ?

mathslover (mathslover):

Let \(f(x) = \int_{2}^{x} \frac{dt}{\sqrt{1+t^4}} \) and \(g\) be the the inverse of \(f\). Then, the value of \(g'(0)\) is ... ?

OpenStudy (rational):

Hint : (2, 0) is a point on f(x)

mathslover (mathslover):

I wished to follow the conventional way. That is to first find \(f\) and then find it's inverse and carry on this way. Here is what I did till now: I did a substitution of \(t^2 = \tan \theta\) which gives : \(2t dt = \sec ^2 \theta \implies dt = \cfrac{\sec^2 \theta}{\sqrt{\tan \theta}} \) The integral becomes: \[\int_{2}^{x} \cfrac{\sec \theta}{2\sqrt{\tan \theta}} d \theta \] Should I follow the product rule or another substitution?

OpenStudy (rational):

clearly thats very painful, we don't need to find g(x), the question is simply about finding g'(0)

mathslover (mathslover):

Can you just show me how does that hint help here? And how did you arrive to that?

mathslover (mathslover):

My book found \(f'(x)\) first and then found \(g'(x)\). I'm not sure how did it do that though.

OpenStudy (rational):

In light of the earlier hint, start by finding \(f'(x)\) : \[f(x) = \int_{2}^{x} \frac{dt}{\sqrt{1+t^4}} \implies f'(x) = \frac{1}{\sqrt{1+x^4}}\]

mathslover (mathslover):

Yeah, that's what my book shows! But, am very unsure how did you arrive to \(f'(x)\) . Why didn't we consider the lower limit 2 ? Why only x?

OpenStudy (rational):

lets work that in steps

OpenStudy (rational):

\[f(x) = \int_{2}^{x} \frac{dt}{\sqrt{1+t^4}} \] let \(F(t) = \int \frac{dt}{\sqrt{1+t^4}}\implies F'(t) = \frac{1}{\sqrt{1+t^4}} \) yes ?

mathslover (mathslover):

Yeah!

OpenStudy (rational):

Next apply fundamental theorem of calculus : \[f(x) = \int_{2}^{x} \frac{dt}{\sqrt{1+t^4}} = F(x)-F(2) \] still yes ?

mathslover (mathslover):

Ahh.. I see, and derivative of \(F(2)\) will equal zero.

OpenStudy (rational):

Exactly!

mathslover (mathslover):

That's awesome. So, next we will find \(g'(x)\) . Just a second, I'll give it a try.

OpenStudy (rational):

k

mathslover (mathslover):

 What does this mean?